For a standard normal distribution, what is the probability that z is greater than 1.96?

A. 0.9750

B. 0.0250

C. 0.0500

D. .5025

E. 0.4750
Answer #1

Solution: From normal distribution table; P (z > 1.96) Hence, option-B is the correct answer 1-P(z 1.96) = 1-0.9750 = 0.0250